OpenStudy (anonymous):

Find a cubic function with the given zeros. 6, -5, 2 How would I solve this?

OpenStudy (jdoe0001):

\(\large { x=\{6, -5, 2\}\implies \begin{cases} x=6\to &x-6=0\\ x=-5\to &x+5=0\\ x=2\to &x-2=0 \end{cases} \\ \quad \\ (x-6)(x+5)(x-2)=\textit{original polynomial} }\)

OpenStudy (jdoe0001):

3 values for "x", means, 3 factors in this case, and thus a 3rd degree polynomial or namely a "cubic"

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

how would i solve it? go to reverse.